Consider the Keynesian Consumption function

C 2000 + 0.8 (Y - T)

C is personal consumption

Y is personal income

T is income taxes

Also suppose Y = 40,000

T = 0.25Y

Calculate APC and APS

Calculate the Multiplier and interpret its value
Answer #1

C = 2000 + 0.8 (Y - 0.25Y) = 2000 + 0.8(0.75Y) = 2000 + 0.8(0.75*40,000) = 2000 + 0.8(30,000) = 2000 + 24,000 = 26,000

Yd = Y - T = Y - 0.25Y = 0.75Y = 0.75*40,000 = 30,000

APC = C/Yd = 26,000/30,000 = 0.87
So, APC = 0.87

APS = 1 - APC = 1 - 0.87 = 0.13
So, APS = 0.13

MPC = 0.8
Multiplier = 1/(1-MPC) = 1/(1-0.8) = 1/0.2 = 5

This means that if spending increases by 1 unit, income will increase by 5 units.